# D Scaling Up Deepfake Detection by Learning from Discrepancy

### v1

@@ 4.1 Scaled-Up Datasets
The GenImage dataset comprises 2,681,167 images, segregated into 1,331,167 real (drawn from ImageNet) and 1,350,000 fake images, with 50k images left for testing in each generator type. Fake images in this dataset are generated by 8 generators, including 1 GAN (BigGAN [4]) and 7 diffusion models (Stable Diffusion V1.4 [36], Stable Diffusion V1.5 [36], GLIDE [31], VQDM [20], Wukong [3], ADM [13] and Midjourney [2]).
In our new setup, we merge the two datasets into one and regard generators with the same architecture but different parameters (e.g. SD v1.4 and SD v1.5) as the same type. To this end, we eliminate SD v1.5 from the merged training set and remain SD v1.4. Besides, we also make sure each generator has an equal number of samples to avoid imbalanced data. Finally, there are a total of 20 generators available in the scale-up setup, 8 in the training set (2 GANs and 6 Diffusions) and 12 in the OOD set. Compared with the former setup, we have scaled up both the number of generators and the volume of samples within the training set. Besides, we further split 10% samples in the testing set of our merged dataset per generator as the validation and the rest as the testing set.
@@ 4.2 Implementation Details
In our experiments, we employ the CLIP: VIT-L/14 model pre-trained on a large dataset WebImageText as the backbone to extract patch tokens. We adopt its penultimate feature before dimensionality reduction for better representation. With the backbone frozen, the whole network is trained on a single NVIDIA RTX 4090 with a batch size of 128 within 5 epochs. We apply the Adam optimizer with an initial learning rate of 0.0001 and a weight decay of 0. During training, all images are resized to 256x256 resolution and then randomly cropped to 224x224 resolution. We adopted the data augmentation scheme the same as the works [45; 32], involving Gaussian blur and JPEG compression. Specifically, the quality of JPEG compression randomly ranges from 30 to 100, the sigma of Gaussian blur randomly ranges from 0 to 3, and their occurrence probabilities are set to 0.5. During validation and testing, we directly resize the image to the size of 224x224, without any augmentation.
4.3 Experimental Setup
Implementation of Baselines. In terms of baselines, we re-implement the comparison methods according to their official codes under our problem settings for fairness. The baselines include CNNDet [45], Patchfor [5], LNP [29], DIRE [46], and UFD [32]. Detailedly, CNNDet finetunes a pre-trained ResNet50 as a classifier and we use the ResNet50 version with default parameters. Patchfor finetunes the truncated ResNet or Xception as the classifier at the patch level and we choose their Xception version. LNP pre-trains a reconstruction model with only real images to extract noise patterns and then finetunes the pre-trained ResNet50 with spatial and spectrum information. We use their pre-trained reconstruction model to preprocess all the images and finetune the ResNet50 in the default way. DIRE first computes the diffusion reconstruction error of all images and trains a ResNet50 on them. We use their 256x256 unconditional DMs to preprocess all the images and also finetune the ResNet50 in the default way. UFD extracts the features with a frozen CLIP model and classifies them by nearest neighbors or linear probing. We select their linear probing version with the CLIP: VIT-L/14 as the backbone. In our experiments, all methods are re-implemented with their default settings in our new setup, and the optimal checkpoint is selected based on the performance on the validation set to ensure a fair comparison.
AP quantifies the separability between fake and real images, neglecting the specific threshold consideration. Previous literature [32; 29; 46; 45] typically uses mAP to represent the overall AP score, where they calculate AP for each generator independently and average the AP scores. However, the direct average of AP disregards the substantial variance in the optimal thresholds, or boundaries, between real and fake images from different generators. In practice, it is infeasible for users to set different real/fake thresholds for unseen generators. Therefore, we propose to calculate a global AP rather than mAP, which shares the same standards among all testing generators.
4.4 Scaling Up the Training Data
In this section, we conduct experiments to show how the models perform as the generators gradually scale up in Fig. 2. We re-train the detectors on the training set with {1, 2, 4, 8} generators. To mitigate any potential bias, we add the generator in three distinct random orders and average the corresponding test results. To make all the curves comparable (in both horizontal and vertical), all detectors are evaluated on the same ID and OOD test sets, thus the curves are comparable in both horizontal and vertical and we can observe how a model baheves as the number of training generators grows. Note that only the first generator is commonly shared when gradually adding training generators from one to eight, we thus test all models only on this generator for the ID evaluation. For the OOD evaluation, we use the 12 common unseen generators for all models.
According to Fig. 2, as the number of generators increases, our method shows a slight fluctuation in ID performance and exhibits the highest performance and stable upward trend in OOD generalization. CNNDet [45] and Patchfor [5] show excellent ID fitting ability, but their OOD performances are inferior to the other works. It is worth mentioning that Patchfor to some extent alleviates the overfitting issue by focusing on artifacts of local patches instead of the high-level features. Reconstruction-based methods like LNP and DIRE heavily rely on their pre-trained reconstruction model, which could result in poor generalization to unfamiliar domains. Moreover, the CLIP-based method UFD exhibits superior OOD performance when there are few generators and maintains a rising trend as the training set scales up. However, its ID performance continues to decline, which indicates a limitation in its representative capacity. In sum, these findings collectively demonstrate that our method effectively improves the distinguishability and generalizability of features by patch-shuffling and artifact invariance learning from the introduced discrepancy, resulting in stable and outstanding performance of ID fitting and OOD generalization.
| Methods | Pub. | Testing | |||||
| In-domain | Out-of-domain | Total | |||||
@@ 4.5 Comparison with the State-of-the-Art Methods
We compare our method with other state-of-the-art methods in the scale-up setup and report evaluation of their ID and OOD performance on the testing set. Tab. 1 reports the results on the testing set, where “In-domain” includes 8 types of generators, and “Out-of-domain” includes 12 types of generators, and “Total” includes all of them. A more detailed version of the result can be found in the Appendix 0.D.
@@
The experimental results exhibit the strengths and weaknesses of each method. CNNDet overfits data from familiar generators and performs poorly on data from unseen generators, consistent with the findings about ResNet in the paper [52]. Patchfor achieves the best ID mean accuracy and AP. However, its subpar OOD performance suggests that directly learning the specific local artifact patterns may also lead to a higher potential for overfitting. LNP exhibits poor performance in both ID and OOD fitting. DIRE demonstrates superior ID performance, surpassed only by Patchfor. However, it lacks generalization capabilities with OOD generators due to its tailored design for DMs. UFD’s performance is second only to our method, but its ID performance is subpar.
@@
In comparison, our method achieves the highest overall mean accuracy of 90.7% and AP of 96.8%, exceeding the second-best method, Patchfor, in overall mean accuracy by 4.2% and AP by 2.6%. Although our method marginally trails Patchfor in terms of ID accuracy and AP, it exhibits a more balanced and consistent overall performance. Furthermore, our method greatly outperforms UFD, which also leverages the pre-training CLIP, in ID mean accuracy by 10% and OOD mean accuracy by 5.3%. This underscores that our method further exploits the potential of CLIP-based feature representation ability by learning from the discrepancy of the shared artifact in the original image and its patch-shuffled variant.
@@
Robustness Evaluation. We evaluate the robustness of all methods to align more closely with their practical effectiveness in real-world applications. Given the potential for real-world images to undergo various degradations, ensuring the robustness of detectors against unseen perturbations becomes another vital aspect of evaluating their performance. A detector that can resist various perturbations and attacks has higher potential and value for practical application. Therefore, we evaluate all methods on two post-processing operations, Gaussian blurring (sigma: 0-2) and JPEG compression (quality: 30-100). Fig. 3 presents the results of our comparison with the other methods, demonstrating that our method exhibits the best robustness with a large margin with the other methods. Note that except for Patchfor and DIRE, all the other methods employ an augmentation strategy involving random Gaussian blurring and JPEG compression during training to enhance robustness. Despite this, we can observe that only our method and UFD exhibit general robustness, even to extremely heavy degradations. In addition, from the perspective of accuracy and the decay rate, our method shows obvious advantages compared to UFD. Discrepancy learning between the original image and its patch-shuffled one can to some extent offset the impact of these disturbances due to its concentration on more stable and invariant artifacts instead of specific artifact patterns.
4.6 Exploration of Different Disruptions
In this subsection, we investigate how different levels of disruptions influence model performance. We conduct experiments by only replacing the patch-shuffling operation with other disruption operations in our framework. These operations include horizontal flipping, vertical flipping, and random rotating (rotating an image with random degrees from 0° to 180°).
The experimental results are shown in Table 2, showing that disrupted feature helps in promoting the model performance, with the effectiveness of patch-shuffling, random rotating, vertical flipping, and horizontal flipping diminishing in sequence. The two-branch model with the horizontal flipping doesn’t show superior to the one-branch baseline, because the horizontally-flipped image could not provide any discrepancy and its features are just the same as the original image for most existing vision models. Comparatively, vertical flipping has provided corruptions to the vision backbones to some degree since most vision models could not well recognize the image after vertical flipping. Thus the discrepancy helps the model learn better artifacts, leading to an overall performance boost of 4% on mean accuracy compared to the horizontal flipping. In addition, random rotation randomly changes the angle of images, introducing a slightly stronger semantic corruption on image semantics than the fixed 180° angle of vertical flipping, and thus has a slightly better overall performance. In comparison to the three operations, the patch-shuffling operation disrupts the image more completely, leading to the maximum discrepancy between the two branches. Thus it achieves the best performance. The comparison between different transformations clearly illustrates that the discrepancy between the original images and corrupted images helps the detector to learn more universal artifacts. The stronger semantic corruption the operation causes, the better the model performs.
| Augmentations | Validataion | |||||
| In-domain | Out-of-domain | Total | ||||
| Mean acc. | AP | Mean acc. | AP | Mean acc. | AP | |
| Horizontal flipping | 0.887 | 0.964 | 0.819 | 0.934 | 0.846 | 0.945 |
| Vertical flipping | 0.934 | 0.986 | 0.854 | 0.937 | 0.886 | 0.959 |
| Random rotation | 0.954 | 0.991 | 0.854 | 0.928 | 0.891 | 0.958 |
| Patch shuffling | 0.972 | 0.995 | 0.874 | 0.931 | 0.913 | 0.962 |
4.7 Ablation Studies
In this subsection, we break down our framework into several key modules and explore the role of each module through ablation experiments. The ablation experimental results are shown in Table 3, where a tick represents the adoption of this module and the emptiness of Branch2 means using only one branch. The experiments have been numbered with groups for the following reference.
In conclusion, each module is dispensable. The impact of any individual module may not be substantial, but their collective integration significantly enhances the performance. These findings provide robust evidence in support of our proposition that invariance-learning from discrepancy is beneficial for universal deepfake detection in a scaled-up scenario.
| Group | Branch1 | Branch2 | CLIP | SA | FC | Validataion | |||||
| In-domain | Out-of-domain | Average | |||||||||
@@ 5 Conclusion and Discussion
Conclusion. In this paper, we propose a novel evaluation setup “train-on-one and test-on-many”, aimed at advancing the development of a universal deepfake detection system. To tackle the challenges in the new setup, we introduce our Discrepancy Deepfake Detector (D) to learn the universal artifacts from multiple generators. We achieve this by introducing a parallel network branch that takes a distorted image as an extra discrepancy signal to supplement its original counterpart. Extensive experiments demonstrate the effectiveness of our framework, achieving 5.3% accuracy improvement in the OOD testing compared to the current SOTA methods and maintaining excellent ID performance.
Limitations. The parallel design of D requires running the CLIP twice to obtain two different features from one image, which doubles the computation cost compared with UFD. Moreover, our proposed setup demands a large-scale and comprehensive real-world dataset, while the merged dataset in our research may not fully meet this criterion in the future. Collecting such a high-quality generated image detection dataset may be expensive, however, it is not infeasible given the fast-increasing research scale in the community.
Appendix 0.A Sensitivity of Discrepancy Features to Patch Size
@@
We study how the patch size affects the learning from the discrepancy features. A tradeoff exists between increasing the discrepancy between features and mining the universal local artifacts, i.e. the smaller patch size offers more discrepancy but retains fewer local artifacts useful for deepfake detection. Therefore, given the original image size of 224, we conduct validations with different patch sizes, ranging from 1 to 224, to see the changing trend. These experiments adhere to the proposed experimental setting, with the patch size being the only variable adjusted. As shown in Tab. 4, changing the patch size from 224 to 112 brings a significant improvement of 6.0 points in ID accuracy and 2.9 points in OOD accuracy, suggesting that introducing additional discrepancy in features helps in expanding the representation of features and extracting universal artifacts.
The patch sizes 14, 28, and 56 yield similarly high overall performance, showing the robustness of introducing discrepancy in different patch sizes. It is worth noting that when the patch size is 1, the local artifacts of the patch-shuffled image are significantly affected, resulting in a drop in model performance compared to patch size 14. In our state-of-the-art version, we directly opt for a patch size of 14 to align with the backbone CLIP:ViT-L/14 [34] while introducing the highest discrepancy in features.
Appendix 0.B Different Classifier Heads

### v2

@@ 4.1 Scaled-Up Datasets
The GenImage dataset comprises 2,681,167 images, segregated into 1,331,167 real and 1,350,000 fake images, with 50k images left for testing in each generator type. Fake images in this dataset are generated by 8 generators, including 1 GAN (BigGAN [4]) and 7 diffusion models (Stable Diffusion V1.4 [34], Stable Diffusion V1.5, GLIDE [29], VQDM [18], Wukong [3], ADM [13] and Midjourney [2]).
In our new setup, we merge two datasets into one and treat generators with the same architecture but different parameters (e.g. SD v1.4 and SD v1.5) as the same type. To this end, we eliminate SD v1.5 from the merged training set and remain SD v1.4. Besides, we make sure each generator has an equal number of samples to avoid imbalanced data. Finally, there are 20 generators available in the scale-up setup, 8 in the training set (2 GANs and 6 Diffusions) and 12 in the OOD set. We further split 10% samples per generator of the testing set as the validation set.
@@ 4.2 Implementation Details
In our experiments, we employ the CLIP: VIT-L/14 model pre-trained on a large dataset WebImageText as the backbone to extract patch tokens. We adopt its penultimate feature before dimensionality reduction for better representation. With the backbone frozen, the whole network is trained on a single NVIDIA RTX 4090 with a batch size of 128 within 5 epochs. We apply the Adam optimizer with an initial learning rate of 0.0001 and a weight decay of 0. During training, all images are resized to 256x256 resolution and then randomly cropped to 224x224 resolution. We adopted the data augmentation scheme the same as the works [42, 30], involving Gaussian blur and JPEG compression. Specifically, the quality of JPEG compression randomly ranges from 30 to 100, the sigma of Gaussian blur randomly ranges from 0 to 3, and their occurrence probabilities are set to 0.5. During validation and testing, we resize the image to size 224x224, without any augmentation.
| Methods | Pub. | In-domain | Out-of-domain | Total | |||
@@
4.3 Experimental Setup
Implementation of Baselines. In terms of baselines, we re-train the comparison methods according to their official codes under our problem settings for fairness. The baselines include CNNDet [42] (ResNet50 version), Patchfor [5] (Xception version), LNP [27], DIRE [43], and UFD [30] (linear probing version). In our experiments, all methods are re-implemented with their default settings in our new setup, and the optimal checkpoint is selected based on the validation set performance to ensure a fair comparison.
AP quantifies the separability between fake and real images, neglecting the specific threshold consideration. Previous literature [30, 27, 43, 42] typically uses mAP to represent the overall AP score, where they calculate AP for each generator independently and average the AP scores. However, the direct average of AP disregards the substantial variance in the optimal thresholds, or boundaries, between real and fake images from different generators. In practice, it is infeasible for users to set different real/fake thresholds for unseen generators. Therefore, we propose to calculate a global AP rather than mAP, which shares the same standards among all testing generators.
4.4 Scaling Up the Training Data
In this section, we conduct experiments to show how the models perform as the generators gradually scale up in Fig. 2. We re-train the detectors on the training set with {1, 2, 4, 8} generators. To mitigate any potential bias, we add the generator in three distinct random orders and average the corresponding test results. To make all the curves comparable (in both horizontal and vertical), all detectors are evaluated on the same ID and OOD test sets, thus the curves are comparable in both horizontal and vertical and we can observe how a model behaves as the number of training generators grows. Note that only the first generator is commonly shared when gradually adding training generators from one to eight, we thus test all models only on this generator for the ID evaluation. For the OOD evaluation, we use the 12 common unseen generators for all models.
According to Fig. 2, as the number of generators increases, our method shows a slight fluctuation in ID performance and exhibits the highest performance and stable upward trend in OOD generalization. CNNDet [42] and Patchfor [5] show excellent ID fitting ability, but their OOD performances are inferior to the other works. It is worth mentioning that Patchfor to some extent alleviates the overfitting issue by focusing on artifacts of local patches instead of the high-level features. Reconstruction-based methods like LNP and DIRE heavily rely on their pre-trained reconstruction model, which could result in poor generalization to unfamiliar domains. Moreover, the pretrained-based method UFD exhibits superior OOD performance when there are few generators and maintains a rising trend as the training set scales up. However, its ID performance continues to decline, which indicates a limitation in its representative capacity. These findings demonstrate that effectively improves the distinguishability and generalizability of features by patch-shuffling and artifact invariance learning from the introduced discrepancy, resulting in outstanding performance of ID fitting and OOD generalization.
@@ 4.5 Comparison with the State-of-the-Art Methods
We compare our method with other state-of-the-art methods in the scale-up setup and report an evaluation of their ID and OOD performance on the testing set. Tab. 1 reports the results on the testing set, where “In-domain” includes 8 types of generators, and “Out-of-domain” includes 12 types of generators, and “Total” includes all of them. A more detailed version of the result can be found in the appendix.
@@
The experimental results exhibit the strengths and weaknesses of each method after scaled-up. CNNDet overfits data from familiar generators and performs poorly on data from unseen generators, consistent with the findings about ResNet in the paper [50]. Patchfor achieves the best ID mean accuracy and AP. However, its subpar OOD performance suggests that directly learning the specific local artifact patterns may also lead to a higher potential for overfitting. LNP exhibits poor performance in both ID and OOD fitting. DIRE demonstrates superior ID performance, surpassed only by Patchfor. However, it lacks generalization capabilities with OOD generators due to its tailored design for DMs. UFD ranks second only to our method. However, its ID performance is significantly inferior.
@@
D achieve the best overall performance under a fair comparison. In comparison, our method achieves the highest overall mean accuracy of 90.7% and AP of 96.8%, exceeding the second-best method, Patchfor, in overall mean accuracy by 4.2% and AP by 2.6%. Although our method marginally trails Patchfor in terms of ID accuracy and AP, it exhibits a more balanced and consistent overall performance. Furthermore, our method greatly outperforms UFD, which also leverages the same pre-training vision backbone, in ID mean accuracy by 10% and OOD mean accuracy by 5.3%. This underscores that our method further exploits the potential of pretrained-based feature representation ability by learning from the discrepancy of the shared artifact in the original image and its patch-shuffled variant.
@@
Robustness Evaluation. We evaluate the robustness of all methods to align more closely with their practical effectiveness in real-world applications. Given the potential for real-world images to undergo various degradations, ensuring the robustness of detectors against unseen perturbations becomes another vital aspect of evaluating their performance. A detector that can resist various perturbations and attacks has higher potential and value for practical application. Therefore, we evaluate all methods on two post-processing operations, Gaussian blurring (sigma: 0-2) and JPEG compression (quality: 30-100). Fig. 3 presents the results of our comparison with the other methods, demonstrating that our method exhibits the best robustness with a large margin with the other methods. Note that except for Patchfor and DIRE, all the other methods employ an augmentation strategy involving random Gaussian blurring and JPEG compression during training to enhance robustness. Despite this, we can observe that only our method and UFD exhibit general robustness, even to extremely heavy degradations. In addition, from the perspective of accuracy and the decay rate, our method shows obvious advantages compared to UFD. Discrepancy learning between the original image and its patch-shuffled one offsets the impact of these disturbances due to its concentration on more stable and invariant artifacts instead of specific artifact patterns.
| Group | Branch1 | Branch2 | CLIP | SA | FC | In-domain | Out-of-domain | Average | |||
@@
4.6 Exploration of Different Disruptions
In this subsection, we investigate how different levels of disruptions influence model performance. We conduct experiments by only replacing the patch-shuffling operation with other disruption operations in our framework. These operations include horizontal flipping, vertical flipping, and random rotating (randomly rotating an image from 0° to 180°).
The experimental results shown inTab. 3, show that each disrupted feature helps in promoting the model performance, with the effectiveness of patch-shuffling, random rotating, vertical flipping, and horizontal flipping diminishing in sequence. The two-branch model with the horizontal flipping doesn’t show superior to the one-branch baseline, because the horizontally-flipped image could not provide any discrepancy and its features are just the same as the original image for most existing vision models. Comparatively, vertical flipping has provided corruptions to the vision backbones to some degree since most vision models could not well recognize the image after vertical flipping. Thus the discrepancy helps the model learn better artifacts, leading to an overall performance boost of 4% on mean accuracy compared to the horizontal flipping. In addition, random rotation randomly changes the angle of images, introducing a slightly stronger semantic corruption on image semantics than the fixed 180° angle of vertical flipping, and thus has a slightly better overall performance. In comparison to the three operations, the patch-shuffling operation disrupts the high-level image semantics more completely while preserving low-level information, leading to an informative discrepancy between the two branches. Thus it achieves the best performance. The comparison between different transformations illustrates that the discrepancy between the original images and corrupted images helps the detector learn more universal artifacts. The stronger the corruption the operation causes, the better the model performs.
| Augmentations | In-domain | Out-of-domain | Total | |||
| Mean acc. | AP | Mean acc. | AP | Mean acc. | AP | |
| Horizontal flipping | 0.887 | 0.964 | 0.819 | 0.934 | 0.846 | 0.945 |
| Vertical flipping | 0.934 | 0.986 | 0.854 | 0.937 | 0.886 | 0.959 |
| Random rotation | 0.954 | 0.991 | 0.854 | 0.928 | 0.891 | 0.958 |
| Patch shuffling | 0.972 | 0.995 | 0.874 | 0.931 | 0.913 | 0.962 |
4.7 Ablation Studies
In this subsection, we break down our framework into several key modules and explore the role of each module through ablation experiments. The ablation experimental results are shown in Tab. 2, where a tick represents the adoption of this module and the emptiness of Branch2 means using only one branch. The experiments have been numbered with groups for the following reference.
In conclusion, each module is dispensable. The impact of any module may not be substantial, but their collective integration significantly enhances the performance. These findings provide robust evidence to support our proposition that learning from discrepancy is beneficial for universal deepfake detection in a scaled-up scenario.
@@ 5 Conclusion and Discussion
Conclusion. In this paper, we propose a novel evaluation setup “train-on-many and test-on-many”, aimed at advancing the development of a universal deepfake detection system. To tackle the challenges in the new setup, we introduce our Discrepancy Deepfake Detector (D) to learn the universal artifacts from multiple generators. We achieve this by introducing a parallel network branch that takes a distorted image as an extra discrepancy signal to supplement its original counterpart. Extensive experiments demonstrate the effectiveness of D, achieving 5.3% accuracy improvement in the OOD testing compared to the current SOTA methods and maintaining excellent ID performance.
Limitations. The parallel design of D requires running the vision model twice to obtain two different features from one image, which doubles the computation cost compared with UFD. Our proposed setup demands a large-scale and comprehensive dataset, while our merged dataset may not fully meet this criterion in the future. Collecting such a high-quality detection dataset may be expensive, however, not infeasible given the fast-increasing research scale.
6 Acknowledgements
This work was supported by the National Natural Science Foundation of China under grant 62372341. Research reported in this was partially supported by an Amazon Research Award Spring 2024 to OR and YZ. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Amazon.
Supplementary Material
In the appendix, we add some experiments to delve deeper into the robustness of the differentiated features. We conduct a sensitivity analysis on the only variable of the patch-shuffle operation, the patch size, in Sec. A; we use different classification heads to demonstrate the robustness of the differentiated features to the classification heads in Sec. B; we showcase the samples that were correctly identified additionally by using differentiated features compared to the baseline UFD in Sec. C; we detailly report the test results of our method and the baseline across various generators in Sec. D; finally, we visualize the focus of UFD and to demonstrate ’s ability of capturing more comprehensive and universal artifacts in Sec. E.
Appendix A Sensitivity of Discrepancy Features to Patch Size
@@
We study how the patch size affects the learning of discrepancy signals. A tradeoff exists between increasing the discrepancy between features and mining the universal local artifacts, i.e. the smaller patch size offers more discrepancy but retains fewer local artifacts. Therefore, given the original image size of 224, we conduct validations with different patch sizes, ranging from 1 to 224, to see the changing trend. These experiments adhere to the previous setting, with only patch size being adjusted. As shown in Tab. 1, changing the patch size from 224 to 112 brings a significant improvement of 6.0 points in ID accuracy and 2.9 points in OOD accuracy, suggesting that additional discrepancy in features helps in expanding the representation of features and extracting universal artifacts. The patch sizes 14, 28, and 56 yield similarly high overall performance, showing the introduced discrepancy’s robustness in different patch sizes. Note that when the patch size is 1, the local artifacts of the shuffled image are significantly affected, resulting in a drop in model performance compared to patch size 14. In our SOTA version, we directly opt for a patch size of 14 to align with the backbone CLIP:ViT-L/14 [32] while introducing the highest discrepancy in features.
Appendix B Different Classifier Heads